OpenStudy (anonymous):

Help! MEDAL AND FAN! A pair of shoes usually sells for $69. If the shoes are 30% off, and sales tax is 6%, what is the total price of the shoes, including tax? $51.68 $47.54 $54.86 $51.20

OpenStudy (anonymous):

so how do you think you should approach this?

OpenStudy (mathstudent55):

You need to do two things. First, you need to take off 30% from the regular price of $69. Then you need to add 6% to the discounted price.

OpenStudy (mathstudent55):

Do you know how to increase a price by a percentage or decrease it by a percentage?

OpenStudy (anonymous):

you have the initial value of 69$, and it is 30% off, so you want to find 30% of 69 dolars

OpenStudy (anonymous):

the of is a multiplication sign and 30% is .3

OpenStudy (anonymous):

you want to multiply to initial ammount by the percentage as a decimal

OpenStudy (anonymous):

that will give you the discounted ammount so you would subtract that value from 69$

OpenStudy (anonymous):

no answers flowers

OpenStudy (anonymous):

then once you find that discounted price \[69\times.3\], you need tosubtract that from 69. the you have your sale price value without the sales tax.

OpenStudy (anonymous):

@MlpSisters7 you still there

OpenStudy (anonymous):

Sorry os was lagging again.. -.-

OpenStudy (anonymous):

20.7

OpenStudy (anonymous):

yea so thats your value you subtract from 69

OpenStudy (anonymous):

Once you have that value you need to multiply that by .06. and add that value to the discounted price.

OpenStudy (anonymous):

\[part 1) 69\times.3=x...part 2) 69-x=y... part 3)y \times.06=z...part4)y+z=discouted price\]